OpenStudy (anonymous):

hi,a basic question, i added 20 ml of 1M Kno3 with 20 ml of 2 mM of Fe3CN6. will the concentration of the two chemicals remain same in mixture??

OpenStudy (anonymous):

What is the balanced chemical equation for this reaction?

OpenStudy (anonymous):

no idea :(

OpenStudy (anonymous):

And what does M and mM stand for?

OpenStudy (anonymous):

Oh so this was an experiment then?

OpenStudy (anonymous):

M means molar mM means mili molar

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

1 mole= 1000 mili moles, so 1 mili mole = 1/1000 moles. 2 mili moles is therefore 2/1000 moles or 0.002 moles. CKNO3= n/V= 1 mol /0.02 L= 50 mol/L CFe3CN6= n/V = 0.002 mol /0.02 L = 0.1 mol/L This is how I usually do it and my teacher does. Note: Where C is concentration, V is volume (in litres) and n is moles. The concentration is therefore different for both chemicals in solution.

OpenStudy (aaronq):

That's not quite right. For example, you have 20 mL of 1 M KNO3. (Are you sure it's not KMnO3?) M= n/L. -> n=M*L = 1M*0.02 L = 0.02 moles New volume 20mL+ 20mL =40 mL = 0.04 L New molarity of KNO3 = 0.02 moles/0.04 L = 0.5 M

OpenStudy (anonymous):

@aaronq,AnImEfReaK thanks for reply....... i think what aaronq says, makes sense....

OpenStudy (preetha):

What is Fe3CN6?

OpenStudy (anonymous):

sorry......typo.... K3[FeCN6] - potassium ferrocyanide